OpenStudy (anonymous):

Part A: Maria rented a coat at $285 for 3 days. If she rents the same coat for 6 days, she has to pay a total rent of $510. Write an equation in the standard form to represent the total rent (y) that Maria has to pay for renting the coat for x days. (4 points) Part B: Write the equation obtained in Part A using function notation.(2 points) Part C: Describe the steps to graph the equation obtained above on the coordinate axes. Mention the labels on the axes and the intervals. (4 points)

OpenStudy (ganpat):

y = a +bx a is fixed cost, while b is variable cost so equation 1 & 2 are as follows, 285 = a +3b 510 = a + 6b can u solve for a & b ??

OpenStudy (ganpat):

hellloo !! anybody home ??

OpenStudy (ganpat):

Find a & b, substitute back in basic equation i.e y = a +bx.. thats solution for part A ..

OpenStudy (anonymous):

Oh sorry i didnt realize anybody answered my question!!

OpenStudy (ganpat):

soo.. now u do have solution for part 1.. rt ??

OpenStudy (anonymous):

can someone help me with a literature question?

OpenStudy (anonymous):

rt ?

OpenStudy (ganpat):

oops !! sorry... right !!

OpenStudy (anonymous):

285 = a +3b 510 = a + 6b is the top equation im suppose to solve for A and bottom B ?

OpenStudy (ganpat):

try solving them simultaneously.. u will get values for a & b..

OpenStudy (anonymous):

a=3n B=95 right ?

OpenStudy (math&ing001):

 replace b with it's value and solve for a.